OpenStudy (anonymous):

how do I do these??? please dont give answer - Do I have to factor it out? \ State the vertex of the following function: f (x ) = (x - 6)^2+ 8 .

Directrix (directrix):

The vertex is a maximum or minimum value of the parabola. In this case, f (x ) = (x - 6)^2+ 8, for the x-coordinate of the vertex, you want to find the value of x that makes (x - 6)^2+ 8 as small as possible.

OpenStudy (anonymous):

-8 and 6?

OpenStudy (jdoe0001):

notice that your equation is already in vertex form http://www.mathwarehouse.com/geometry/parabola/images/standard-vertex-forms.png so your vertex is at \(\large f (x ) = (x -\color{red}{ 6})^2+ \color{red}{8} \)

OpenStudy (jdoe0001):

\(\large \begin{array}{llll} &x&y\\ f (x ) = (x - &6)^2+ &8 \end{array}\)

Directrix (directrix):

@mocham >>>-8 and 6? No. The x value of 6 does minimize the function. So, you have (6,?) as the vertex of the parabola. Now, evaluate (x - 6)^2+ 8 for x = 6 to get the y-coordinate of the vertex.

OpenStudy (anonymous):

6 and 8?

Directrix (directrix):

Yes, write those as coordinates. (6,8)

OpenStudy (anonymous):

(6,8) my vertex =)

Directrix (directrix):

Correct.

OpenStudy (anonymous):

thank you!

Directrix (directrix):

You are welcome.